An all-equity firm is considering the following projects:

Project Beta IRR
W .64     9.5 %
X .75     10.6
Y 1.31     14.1
Z 1.42     17.2

The T-bill rate is 5.2 percent, and the expected return on the market is 12.2 percent.

a.

Which projects have a higher/lower expected return than the firm’s 12.2 percent cost of capital?


b.

Which projects should be accepted?


c.

Which projects will be incorrectly accepted/rejected or correctly accepted/rejected if the firm's overall cost of capital were used as a hurdle rate?

What determines if a project gets incorrectly accepted/reject?
Answer #1

1.
W=5.2%+0.64*(12.2%-5.2%)=9.6800%

X=5.2%+0.75*(12.2%-5.2%)=10.4500%

Y=5.2%+1.31*(12.2%-5.2%)=14.3700%

Z=5.2%+1.42*(12.2%-5.2%)=15.1400%

Y and Z have higher
W and X have lower

2.
Accept X and Z

3.
Incorrectly Accepted Y
Incorrectly Rejected X
Correctly Accepted Z
Correctly Rejected W and Y

4.
Risk